OpenStudy (anonymous):

triangle is 8 inch shorter than the side of a square. perimeter of the square is 46 inches more than the perimeter of the triangle. find the lengthof a side of the square

OpenStudy (anonymous):

squares have equal length sides so 46/4 is the length of a side of the square = 11.5inches

OpenStudy (anonymous):

first lets denote those things: let s be the measue of the side of the square let Pt be the perimeter of the triangle let Ps be the perimeter of the square. by saying that the triangle is 8 inch shorter than the side of a square, you mean perimeter of the triangle is 8 inch shorter than the side of a square so: Pt=s-8 now you also stated that the perimeter of the square is 46 inches more than the perimeter of the triangle which we can mathematically write as: Ps=Pt+46 substituting s-8 to Pt Ps=s-8+46 We know that the perimeter of a square is 4 times its side or simply Ps=4s. Substituting this.... 4s=s-8+46 Now let's solve for s 4s=s-8+46 3s=38 s=38/3 s=12.666666666666666666666666666....inches rounding off will give us s=12.67 inches